# Ideals, Gröbner Bases, and PCPs

### v1

@@ 7 The PCP Theorem with One Composition
Assume that there exist $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are robust assignment-testers with parameters $(R_{1},s_{1},\ell_{1},\delta_{1},\varepsilon_{1},w_{1},\rho_{1})$ and $(R_{2},s_{2},\ell_{2},\delta_{2},\varepsilon_{2},w_{2},\rho_{2})$ respectively. If $\rho_{1}\geq\delta_{2}$, then there also exists a robust assignment-tester $\mathcal{A}_{3}$ with parameters $(R_{3},s_{3},\ell_{3},\delta_{3},\varepsilon_{3},w_{3},\rho_{3})$ where
  • $R_{3}(n)=R_{1}(n)\cdot R_{2}(s_{1}(n)),$
  • $s_{3}(n)=s_{2}(s_{1}(n)),$
  • $\ell_{3}(n)=\ell_{2}(s_{1}(n)),$
  • $\delta_{3}(n)=\delta_{1}(n),$
  • $\varepsilon_{3}(n)=\varepsilon_{1}(n)+\varepsilon_{2}(s_{1}(n))-\varepsilon_{1}(n)\cdot\varepsilon_{2}(s_{1}(n)),$
  • $w_{3}(n)=w_{2}(n),$ [^6] This is not explicitly stated in [DinurReingold] but it follows trivially from the proof.
  • $\rho_{3}(n)=\rho_{2}(n).$
@@
The following is a basic property of standard reductions form Circuit-SAT to $3$-$\mathsf{COLOR}$. One can prove it e.g. by using the standard Tseitin transformation (reducing Circuit-SAT to $3$-SAT) followed by the reduction from $3$-SAT to $3$-$\mathsf{COLOR}$ in [GareyJohnsonStockmeyer, Theorem 2.1].
@@
Lemma 7.5 (Robustization: Lemma 3.6 in [DinurReingold]).
There is an absolute constant $c_{1}$ such that if there is an assignment-tester $\mathcal{A}$ with parameters $(R,s,\ell,\delta,\varepsilon,w)$, then there is also a robust assignment-tester $\mathcal{A}^{\prime}$ with parameters $(R^{\prime}=R,s^{\prime}=c_{1}\cdot s,\ell^{\prime}=\ell,\delta^{\prime}=\delta,\varepsilon^{\prime}=\varepsilon,w^{\prime}=\operatorname{poly}(w),\rho^{\prime}=1/(c_{1}\cdot\ell))$.
@@
There exist absolute constants $c,c_{1}$ such that the following holds. Assume that for every $n\geq 1$, there exist $q=q(n),m=m(n),d=d(n),k=k(n)$ such that $q\geq cd^{3}$ and a variety $V_{n}\subseteq\mathbb{F}_{q}^{m}$ of size $\omega(n)$ constructible in time $\operatorname{poly}(n)$ with extension degree $d$ and Gröbner complexity $k$. Then, for any $\delta=\delta(n)>0,$ there is a robust assignment-tester with parameters
$\left(q^{O(k+(m/\delta))},\operatorname{poly}((d\log q)/\delta),\ell=O(1/\delta),\delta,1/4,q^{O(d)},\frac{1}{c_{1}\cdot\ell}\right)$
@@
The constants $c$ and $c_{1}$ are chosen from Theorem 2.3 and Lemma 7.5 respectively. We can assume without loss of generality that $c\geq 100.$
@@
We now consider the algorithm Tester specified in Algorithm 6 with $c_{2}$ and $c_{3}$ being absolute constants that we will choose below. For any choice of these constants, the algorithm uses $r=O((k+(m/\delta))\log q)$ many random bits, makes $O(1/\delta)$ queries to its input oracles $\sigma$ and $\tau$ where the former is defined over the Boolean alphabet and the latter is defined over an alphabet of size $O(d\log q).$ The reduction $\mathcal{A}$ iterates over all sequences $b$ of $r$ random coin tosses used by the algorithm and for each it produces a Boolean circuit $\psi_{b}$ that performs the checks specified in the algorithm.
1 Input: Degree parameter $d$ and oracle access to $(\sigma,\tau)$ where $\sigma:X\rightarrow\{0,1\}$, $\tau:Y\rightarrow\Sigma$, where $\Sigma$ is the alphabet of the PCP verifier from Theorem 2.3 and $|Y|=q^{O(k+m)}$ is the length of the proof. 2 Run the PCP verifier from Theorem 2.3 on $\tau$ independently $c_{2}$ times. If the PCP verifier rejects even once, then return Reject // Random bits $r^{\prime}=O(c_{2}(k+m)\log q)$, number of queries $\ell^{\prime}=O(1)$, alphabet size $O(d\log q)$. 3 4 Run $\mathcal{LC}_{d}^{\hat{\chi},\hat{\chi}^{\prime}}(v_{0})$ and if the algorithm rejects or returns anything other than 1-1, then return Reject // Random bits $O(m\log q)$, number of queries $O(1)$. 5 62mm Repeat $c_{3}/\delta$ times: Sample random $v\in V_{0}$ and run $\mathcal{LC}_{d}^{\hat{\chi},\hat{\chi}^{\prime}}(v)$. If the algorithm rejects or returns anything other than $\sigma(v)$, then return Reject // Random bits $O((c_{3}/\delta)\cdot m\log q)$, number of queries $O(c_{3}/\delta)$. 7 82mm return Accept Algorithm 6 Tester
The number of circuits produced is $2^{r}=q^{O(k+(m/\delta))}$, each of which queries $\ell=O(1/\delta)$ locations in the string $\sigma\cup\tau.$ The computations performed by $\psi_{b}$ are all efficiently computable functions of the queried bits, and hence the circuit $\psi_{b}$ has size that is polynomial in the number of bits queried, which is $O((d\log q)/\delta).$ The alphabet $\Sigma$ is the same as that of the PCP verifier and hence has cardinality $q^{O(d)}.$
To show that this is an assignment tester, it suffices to argue the completeness and soundness criteria. Completeness is trivial from the definition of the PCP and the properties of the NP-completeness reduction argued above.
For soundness, let us assume that $\sigma:X\rightarrow\{0,1\}$ is $\delta$-far from any satisfying assignment to $\varphi.$ We need to show that the probability that Algorithm 6 rejects is at least $3/4$, since this also means that at least $3/4$ of the circuits $\psi_{b}$ reject.
@@
Finally, since $\chi$ is a proper $3$-coloring of $G$ and $\chi(v_{0})=-1$, it follows that the restriction of $\chi$ to $V_{0}$ defines a satisfying assignment of $\varphi$. Since $\sigma$ is $\delta$-far from any satisfying assignment of $\varphi$, it follows that for each $v\in V_{0}$ chosen in the next step, the probability that $\sigma$ and $\chi$ differ at $v$ is at least $\delta$. Further, by Theorem 3.7, the probability that $\mathcal{LC}_{d}^{\hat{\chi},\hat{\chi}^{\prime}}(v)$ returns $\chi(v)$ or rejects is at least $3/4$ (as in the previous paragraph). Thus, the chance that this step does not reject is bounded by
$\left(1-\frac{3\delta}{4}\right)^{c_{3}/\delta}\leq 1/4$
as long as $c_{3}$ is a large enough constant. This concludes the proof of the soundness of $\mathcal{A}.$
@@
$(q^{O(k+(m/\delta))},\operatorname{poly}((k+(m/\delta))\log q),\ell=O(1/\delta),\delta,1/4,q^{O(d)}).$
In order to make the assignment-tester robust, we simply apply the Robustization lemma Lemma 7.5. This concludes the proof of Lemma 7.6. ∎
@@
For any $\delta_{1}=\delta_{1}(n)>0$, we have a robust assignment-tester with parameters
$(R_{1}=n^{O(1/\delta_{1})},s_{1}=\operatorname{poly}((\log n)/\delta_{1}),\ell_{1}=O(1/\delta_{1}),\delta_{1},\varepsilon_{1}=1/4,w_{1}=\operatorname{poly}(n),\rho_{1}=1/(c_{1}\cdot\ell_{1})).$
@@
Set $q=\operatorname{poly}(\log n)$ a power of $3,$ $V=H^{m}$ for $H\subseteq\mathbb{F}_{q}$ of size $\log n$, and $m=O(\log n/\log\log n).$ The lemma follows from Lemma 7.6 and Corollary 4.4. ∎
@@
For any $c=c(n),\delta_{2}=\delta_{2}(n)>0$, we have a robust assignment-tester with parameters
$(R_{2}=c^{O(cn^{2/c}/\delta_{2})},s_{2}=\operatorname{poly}(c/\delta_{2}),\ell_{2}=O(1/\delta_{2}),\delta_{2},\varepsilon_{2}=1/4,w_{2}=c^{O(c)},\rho_{2}=1/(c_{1}\cdot\ell_{2})).$
@@
Set $V=(\{0,1\}^{m/c}_{\leq 1})^{c}\subseteq\mathbb{F}_{q}^{m}$ for a suitably large $q=\operatorname{poly}(c)$ and $m=O(n^{1/c}).$ The lemma follows from Lemma 7.6 and Corollary 4.5. ∎
@@
We compose the robust assignment-testers from Lemma 7.7 and Lemma 7.8 above.
Choosing $\delta_{1}$ to be any absolute constant in $(0,1)$, $c$ to be a large enough absolute constant such that $R_{2}(s_{1}(n))=\operatorname{poly}(n)$, and $\delta_{2}\leq\rho_{1}$ another absolute constant, we obtain two robust assignment testers that can be composed using Lemma 7.3. This leads to a robust assignment-tester with parameters $R_{3}=\operatorname{poly}(n),$ $s_{3},\ell_{3},w_{3}\in O(1)$ and $\delta_{3},1-\varepsilon_{3}\in\Omega(1),$ which by 7.2 implies the PCP theorem. ∎
Appendix A Relation between Gröbner Bases and Gröbner Generating Sets
In Section 4 we introduced Gröbner generating sets. We here show that Gröbner bases are Gröbner generating sets, and that Gröbner bases are also well-behaved under Cartesian products.

### v2

@@ 7 The PCP Theorem with One Composition
Suppose there exist robust assignment-testers $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with parameters $(r_{1},\ell_{1},a_{1},\rho_{1})$ and $(r_{2},\ell_{2},a_{2},\rho_{2})$ respectively. Then there also exists a robust assignment-tester $\mathcal{A}_{3}$ with parameters $(r_{3},\ell_{3},a_{3},\rho_{3})$ where
  • $r_{3}(n)=r_{1}(n)+r_{2}(a_{1}(n)),$
  • $\ell_{3}(n)=\ell_{2}(a_{1}(n)),$
  • $a_{3}(n)=a_{2}(a_{1}(n)),$
  • $\rho_{3}(n)=\rho_{1}(n)\cdot\rho_{2}((a_{1}(n))).$
Proof.
Following the high-level idea above, for a given Boolean circuit $\varphi$ of size $n$, the reduction $\mathcal{A}_{3}$ outputs a certain number of Boolean circuits as follows. The reduction $\mathcal{A}_{3}$ first runs $\mathcal{A}_{1}$ to produce Boolean circuits $\psi_{1},\psi_{2},\dots,\psi_{R_{1}}$, where $R_{1}=2^{r_{1}(n)}$, each of size at most $a_{1}(n)$. Now for each $\psi_{i}$, the reduction now runs $\mathcal{A}_{2}$ with input as $\psi_{i}$, to produce circuits $\psi_{i,j}$ for $j=1,2,\dots,R_{2}$, where $R_{2}=2^{r_{2}(a_{1}(n))}$ (since the circuit input to $\mathcal{A}_{2}$ is of size at most $a_{1}(n)$). The final output of $\mathcal{A}_{3}$ will be the circuits $(\psi_{i,j})_{i\in[R_{1}],j\in[R_{2}]}$ (hence, $r_{3}(n)=r_{1}(n)+r_{2}(a_{1}(n))$). We note that the running time of $\mathcal{A}_{3}$ is $\operatorname{poly}(R_{1}R_{2},n)$, the size of the output circuits is at most $a_{2}(a_{1}(n))$, and each output circuit depends on at most $\ell_{2}(a_{1}(n))$ variables.
Let $\Sigma_{1}$ be the alphabet corresponding to auxiliary variables of $\mathcal{A}_{1}$ for input circuits of size $n$ and $\Sigma_{2}$ be the alphabet corresponding to the auxiliary variables of $\mathcal{A}_{2}$ for input circuits of size $a_{1}(n)$. Let $X$ denote the variables of $\varphi$, $Y$ denote the auxiliary variables of $\mathcal{A}_{1}$ for the circuit $\varphi$, and $Y_{i}$ denote the auxiliary variables of $\mathcal{A}_{2}$ for the circuit $\psi_{i}$ (for $i\in[R_{1}]$). Thus, the auxiliary variables of $\mathcal{A}_{3}$ are $Y\cup\cup_{i\in[R_{1}]}Y_{i}$ with alphabet $\Sigma_{2}$; i.e., we will treat the bits corresponding to $Y$ as being over $\Sigma_{2}$ as well (this can be without increasing the size and the number of variables each final circuit depends on). We will now show completeness and soundness.
If $\varphi(\sigma)=1$ for some $\sigma:X\to\{0,1\}$, by the completeness of $\mathcal{A}_{1}$, we note that there exists $\tau:Y\to\Sigma_{1}$ such that $\psi_{i}(\alpha_{i})=1$ for all $i\in[R_{1}]$, where $\alpha_{i}:=(\sigma\cup\tau)|_{\textrm{Vars}(\psi_{i})}$. Now, by the completeness of $\mathcal{A}_{2}$, this means that there exists $\tau_{i}:Y_{i}\to\Sigma_{2}$ such that $\psi_{i,j}(\beta_{i,j})$ for all $j\in[R_{2}]$, where $\beta_{i,j}:=(\alpha_{i}\cap\tau_{i})|_{\mathrm{Vars}}(\psi_{i,j})$. Therefore, we see that for the assignment $\sigma\cup\tau\cup_{i\in[R_{1}]}\tau_{i}$, all the circuits $\psi_{i,j}$ are satisfied.
To show soundness, let $\delta:=\delta(\sigma,\varphi^{-1}(1))$ and $\tau:Y\to\Sigma_{1}$ and $\tau_{i}:Y_{i}\to\Sigma_{2}$ be arbitrary assignments to the auxiliary variables of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for $i\in[R_{1}]$. Again, let $\alpha_{i}:=(\sigma\cup\tau)|_{\textrm{Vars}(\psi_{i})}$ and $\beta_{i,j}:=(\alpha_{i}\cap\tau_{i})|_{\mathrm{Vars}}(\psi_{i,j})$. By the robust soundness condition of $\mathcal{A}_{1}$, we have that
$\operatorname*{\mathbb{E}}_{i\in[R_{1}]}[\delta(\alpha_{i},\psi_{i}^{-1}(1))]\geq\rho_{1}\delta.$
Similarly, for each $i\in[R_{1}]$, by the robust soundness of $\mathcal{A}_{2}$, we have that
$\operatorname*{\mathbb{E}}_{j\in[R_{2}]}[\delta(\beta_{i,j},\psi_{i,j}^{-1}(1))]\geq\rho_{2}\delta(\alpha_{i},\psi_{i}^{-1}(1)),$
where $\rho_{2}=\rho_{2}(a_{1}(n))$ since the size of $\psi_{i}$ is at most $a_{1}(n)$. Taking expectation over a uniform random $i\in[R_{1}]$ on both sides of the above equation, we get the desired robust soundness for $\mathcal{A}_{3}$:
$\operatorname*{\mathbb{E}}_{i\in[R_{1}],j\in[R_{2}]}[\delta(\beta_{i,j},\psi_{i,j}^{-1}(1))]\geq\rho_{2}\operatorname*{\mathbb{E}}_{i\in[R_{1}]}[\delta(\alpha_{i},\psi_{i}^{-1}(1))]\geq\rho_{1}\rho_{2}\delta.$
This finishes the proof of the composition lemma. ∎
@@
The following is a basic property of standard reductions form Circuit-SAT to $3$-$\mathsf{COLOR}$. One can prove it e.g. by using the standard Tseitin transformation (reducing Circuit-SAT to $3$-SAT) followed by the reduction from $3$-SAT to $3$-$\mathsf{COLOR}$ in [11, Theorem 2.1].
@@
Lemma 7.5 (Robustization: Lemma 3.6 in [9]).
There is an absolute constant $c_{1}$ such that if there is an assignment-tester $\mathcal{A}$ with parameters $(r,\ell,a,\rho)$, then there is also a robust assignment-tester $\mathcal{A}^{\prime}$ with parameters $(r^{\prime}=r,\ell^{\prime}=c_{1}\ell a,a^{\prime}=c_{1}\ell a,\rho^{\prime}=\rho/(c_{1}\ell))$.
@@
There exist absolute constants $c,\rho>0$ such that the following holds. Assume that for every $n\geq 1$, there exist $q=q(n),m=m(n),d=d(n),k=k(n)$ such that $q\geq cd^{3}$ and a variety $V_{n}\subseteq\mathbb{F}_{q}^{m}$ of size $\omega(n)$ constructible in time $\operatorname{poly}(n)$ with extension degree $d$ and Macaulay complexity $k$. Then, there is a robust assignment-tester with parameters
$\left(r={O((k+m)\log q)},\ell=O(d\log q),a=O(d\log q),\rho\right)$
@@
The constant $c$ is chosen from Theorem 2.3. We can assume without loss of generality that $c\geq 100.$
@@
We now consider the assignment-tester specified in Algorithm 6 with $c_{2}$ being an absolute constant that we will choose below. For any choice of this constant, the algorithm uses $r_{1}=O((k+m)\log q)$ many random bits, makes $\ell_{1}=O(1)$ queries to its input oracles $\sigma$ and $\tau$ where the former is defined over the Boolean alphabet and the latter is defined over an alphabet of size $O(d\log q).$ The reduction $\mathcal{A}$ iterates over all sequences $b$ of $r_{1}$ random coin tosses used by the algorithm and for each it produces a Boolean circuit $\psi_{b}$ that performs the checks specified in the algorithm.
1 Input: Degree parameter $d$ and oracle access to $(\sigma,\tau)$ where $\sigma:X\rightarrow\{0,1\}$, $\tau:Y\rightarrow\Sigma$, where $\Sigma$ is the alphabet of the PCP verifier from Theorem 2.3 and $|Y|=q^{O(k+m)}$ is the length of the proof. 2 Run the PCP verifier from Theorem 2.3 on $\tau$ independently $c_{2}$ times. If the PCP verifier rejects even once, then return Reject // Random bits $O(c_{2}(k+m)\log q)$, number of queries $O(1)$, alphabet size $O(d\log q)$. 3 4 Run $\mathcal{LC}_{d}^{\hat{\chi},\hat{\chi}^{\prime}}(v_{0})$ and if the algorithm rejects or returns anything other than 1-1, then return Reject // Random bits $O(m\log q)$, number of queries $O(1)$. 5 62mm Sample $v\in V_{0}$ uniformly at random and run $\mathcal{LC}_{d}^{\hat{\chi},\hat{\chi}^{\prime}}(v)$. If the algorithm rejects or returns anything other than $\sigma(v)$, then return Reject // Random bits $O(m\log q)$, number of queries $O(1)$. 7 82mm return Accept Algorithm 6 Assignment-Tester
The number of circuits produced is $2^{r_{1}}=q^{O(k+m)}$, each of which queries $\ell_{1}=O(1)$ locations in the string $\sigma\cup\tau.$ The computations performed by $\psi_{b}$ are all efficiently computable functions of the queried bits, and hence the circuit $\psi_{b}$ has size that is polynomial in the number of bits queried, which is $O(d\log q).$ The alphabet $\Sigma$ is the same as that of the PCP verifier and hence has cardinality $q^{O(d)}.$
To show that this is an assignment-tester, it suffices to argue the completeness and soundness criteria. Completeness is trivial from the definition of the PCP and the properties of the NP-completeness reduction argued above.
For soundness, let $\sigma:X\rightarrow\{0,1\}$ be an arbitrary assignment and let $\delta:=\delta(\sigma,\varphi^{-1}(1))$. We need to show that the probability that Algorithm 6 rejects is at least $3\delta/4$, since this also means that at least $3\delta/4$ of the circuits $\psi_{b}$ reject.
@@
Finally, since $\chi$ is a proper $3$-coloring of $G$ and $\chi(v_{0})=-1$, it follows that the restriction of $\chi$ to $V_{0}$ defines a satisfying assignment of $\varphi$. Since $\sigma$ is $\delta$-far from any satisfying assignment of $\varphi$, it follows that for a random $v\in V_{0}$ chosen in the next step, the probability that $\sigma$ and $\chi$ differ at $v$ is at least $\delta$. Further, by Theorem 3.7, the probability that $\mathcal{LC}_{d}^{\hat{\chi},\hat{\chi}^{\prime}}(v)$ returns $\chi(v)$ or rejects is at least $3/4$ (as in the previous paragraph). Thus, the chance that this step rejects is at least $3\delta/4$. This concludes the proof of the soundness of $\mathcal{A}.$
@@
$(O(k+m)\log q,\ell_{1}=O(1),O(d\log q),3/4).$
In order to make the assignment-tester robust, we simply apply the Robustization lemma Lemma 7.5. This yields a robust assignment-tester with parameters
$(O(k+m)\log q,O(d\log q),O(d\log q),3/(4c_{1}\ell_{1}))$
for some constant $c_{1}>0$, and concludes the proof of Lemma 7.6. ∎
@@
There exists a robust assignment-tester with parameters
$(r_{1}=O(\log n),\ell_{1}=O(\log n)^{2},a_{1}=O(\log n)^{2},\rho_{1}=\Omega(1)).$
@@
Set $q=(\log n)^{6}$ a power of $3,$ $V=H^{m}$ for $H\subseteq\mathbb{F}_{q}$ of size $\log n$, and $m=O(\log n/\log\log n)$ such that $|V|=\operatorname{poly}(n)$. By Corollary 4.5, we thus have that the extension degree of $V$ is $d=O((\log n)^{2}/\log\log n)$ and Macaulay complexity is $k=O(\log n/\log\log n)$. The lemma then follows by applying Lemma 7.6. ∎
@@
For every $\alpha>0$, there exists a robust assignment-tester with parameters
$(r_{2}=O(n^{\alpha}),\ell_{2}=O(1),a_{2}=O(1),\rho_{2}=\Omega(1)).$
@@
Set $V=(\{0,1\}^{m/c}_{\leq 1})^{c}\subseteq\mathbb{F}_{q}^{m}$ for suitably large constant integers $c\geq 2/\alpha$ (i.e., we are treating $\alpha$ and $c$ as constants) and $q$ which a power of 3, and $m=O(n^{1/c})$ such that $|V|=\operatorname{poly}(n)$. By Corollary 4.6, we thus have that the extension degree of $V$ is $d\leq c$ and Macaulay complexity is $k=O(m^{2})$. Taking $q\geq cd^{3}$ to be a sufficiently large constant, by Lemma 7.6, we obtain a robust assignment-tester with parameters $(O(n^{\alpha}),O(1),O(1),\Omega(1))$
@@
We compose the $(r_{1},\ell_{1},a_{1},\rho_{1})$ robust assignment-testers from Lemma 7.7 with the $(r_{2},\ell_{2},a_{2},\rho_{2})$ robust assignment-tester from Lemma 7.8 corresponding to $\alpha=1/2$, by using Lemma 7.3. This leads to a robust assignment-tester with parameters $(r_{3},\ell_{3},a_{3},\rho_{3})$, where $r_{3}(n)=r_{1}(n)+r_{2}(a_{1}(n))=O(\log n)$, $\ell_{3}=\ell_{2}(a_{1}(n))=O(1),a_{3}=a_{2}(a_{1}(n))=O(1)$, and $\rho_{3}=\rho_{1}\rho_{2}=\Omega(1)$. Combined with 7.2, this implies the PCP theorem. ∎
Acknowledgments
We are thankful to Yassine Ghannane for pointing us to the relevant definitions of Macaulay bases. We are also thankful to the anonymous STOC reviewers for their valuable comments and bringing the work of Khot [14] to our attention.
Appendix A Relation between Gröbner Bases and Macaulay Bases
In Section 4 we introduced Macaulay bases. We here show that all Gröbner bases are Macaulay bases, and that Macaulay bases are also well-behaved under Cartesian products.